OpenStudy (anonymous):

What are the solutions to log3x + log3(x2 + 2) = 1 + 2log3x? x = –2 x = –1 x = 1 x = 2 There is no true solution.

OpenStudy (anonymous):

If this was a MC test and you know how to take the log_3 of numbers, I say just guess and check

OpenStudy (campbell_st):

ok... just a quick tip... you can't take the log of a negative number... \[\log_{3}(-1) ~~doesn't ~exist\] so you can eliminate A and B

OpenStudy (anonymous):

If not, since this is a question with log_3, we have to take 3 and raisee it to each expression since \[\Large a^{\log_{a}x} =x\]

OpenStudy (campbell_st):

so then you are left with using the log laws the 1st law you need \[\log(a) + \log(b) = \log(ab)\] next you need to start the right hand side \[\log_{a}(a) = 1\] also on the right you need \[a \log(x) = \log(x^a)\] so if you use the log laws... you can gte an equation to solve

OpenStudy (anonymous):

thanks guys i figured it out, it was C and D

OpenStudy (anonymous):

\[\Large \log_{3}x + \log_{3}(x^2 + 2) = 1 + 2\log_{3}x\] \[\Huge 3^{\log_{3}x + \log_{3}(x^2 + 2) = 1 + 2\log_{3}x}\]

OpenStudy (anonymous):

\[\Large x+(x^2+2)=3+x^2\]

OpenStudy (anonymous):

x=1

OpenStudy (campbell_st):

well that's not quite right if you apply the log laws it becomes \[\log_{3})x^3 + 2x) = \log_{3}(3x^2)\] now you can equate things \[x^3 + 2x = 3x^2\] and subsequently you can solve for x.

OpenStudy (campbell_st):

which becomes \[x^3 - 3x^2 + 2x = 0~~~or~~~x(x^2 - 3x + 2) = 0\] factor the quadratic and you can find all the solutions but you should ignore x = 0 as a solution as it doesn't exist

OpenStudy (anonymous):

Wait, where did you get an x^3 from?

OpenStudy (anonymous):

If I'm not mistaken, these are log_3s that we are dealing with, and my math is correct then going with that, if the first erm is just an x term, the next is an x^2 which evaluated at x=1 gives me 1. and the other side is just 1+log_3(1^2) which is just 1 as well @campbell_st